OpenStudy (anonymous):

HElllppppp!!!!!!! 6x − 10 + 14 = 22

HanAkoSolo (jamierox4ev3r):

6x+4=22

HanAkoSolo (jamierox4ev3r):

all you have to do is simplify. Can you solve from there?

HanAkoSolo (jamierox4ev3r):

@Mee_Love(813)

OpenStudy (anonymous):

yes Thanks a lot! :)

HanAkoSolo (jamierox4ev3r):

6x+4=22 -4 -4 ------- 6x=18 x=3

HanAkoSolo (jamierox4ev3r):

if you solve, x equals 3

OpenStudy (uri):

6x − 10 + 14 = 22 6x +4=22 subract 4 both the sides 6x +4=22 -4 -4 ====== 6x=18 divide 18 with 6 x=18/6 x=3